OpenStudy (lovelyharmonics):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 3, -13, and 5 + 4i

OpenStudy (lovelyharmonics):

@surjithayer yay you're still on! can you help me? c:

OpenStudy (anonymous):

complex roots always occur in pairs

OpenStudy (anonymous):

(x-3)(x+13)(x-5-4i)(x-5+4i)

OpenStudy (anonymous):

x=5+4i other root is 5-4i

OpenStudy (anonymous):

Multiply and you get your polynomial

OpenStudy (lovelyharmonics):

wait how do i multiply all that? ◑.◑

OpenStudy (anonymous):

\[ (x-3) (x+13) (x-5-4 i) (x-5+4 i)=x^4-98 x^2+800 x-1599 \]

OpenStudy (lovelyharmonics):

but HOW did you get that .-.

OpenStudy (anonymous):

{(x-5)+4i}{(x-5)+4i}=(x-5)^2-(4i)^2

OpenStudy (anonymous):

i^2=-1

OpenStudy (lovelyharmonics):

youre speaking a different language sir

OpenStudy (anonymous):

i am not speaking a different language. (x-5)^2-16i^2=x^2-10x+25+16 =x^2-10x+41 i only solved complex part. now you get (x-3)(x+13)(x^2-10x+41) now you can multiply

OpenStudy (anonymous):

correction i have written wrong above. it should be {x-5+4i}{x-5-4I} ={(x-5)+4i}{(x-5)-4i}

OpenStudy (lovelyharmonics):

so x^3-13x^2+71x-123 and then i multiply it all by x+13?

OpenStudy (lovelyharmonics):

okay i got it ^.^ thank you!

OpenStudy (anonymous):

yw